Terry did all 25 problems on the test and got an 85. Scores are computed by giving 4 points for each right answer and subtracting 1 point for each wrong answer. How many problems did terry get right and wrong?

Answer by Earlsdon(6294)

Let x = the number of problems Terry got right, so he must have got (25-x) wrong and together, these add up to the total of the 25 problems on the test.
Each correct answer got 4 points so he got 4x points less (25-x)(1) for the wrong answers and these add up to the 85 points he scored on the test.
4x-(25-x)(1) = 85
4x-25+x = 85
5x-25 = 85
5x = 110
x = 22 and...
25-x = 25-22 = 3.
Terry got 22 problems right and 3 problems wrong.